OpenStudy (anonymous):

integral of - infinity to infinity of 1/(x^2+4)

OpenStudy (zzr0ck3r):

note that \(\frac{1}{x^2+1}=\frac{1}{4}\frac{1}{(\frac{x}{2})^2+1}\) so \(\int\frac{1}{x^2+4}dx=\frac{1}{4}\int\frac{1}{(\frac{x}{2}+1)}=2*\frac{1}{4}\tan^{-1}(\frac{x}{2})\)

OpenStudy (anonymous):

so we can see it relates to arctan so we make it that?

OpenStudy (zzr0ck3r):

exactly

OpenStudy (zzr0ck3r):

the 2 on the outside is to deal with the chain rule on the x/2 take the derivative of the answer and you will see what I mean.

OpenStudy (anonymous):

ok so if i got a general formula integral of 1/(x^2+a^2)

OpenStudy (anonymous):

we would facotor out 1/a

OpenStudy (zzr0ck3r):

looks like it \(\int\frac{1}{x^2+a^2}dx=\frac{1}{a}\tan^{-1}(\frac{x}{a})\)

OpenStudy (anonymous):

ok sorry but can you explain how you got x/2 when u factored our 1/4 i want to fully understand the algebra!

OpenStudy (zzr0ck3r):

or even better \(\int\frac{1}{x^2+b}dx=\frac{1}{\sqrt{b}}\tan^{-1}(\frac{x}{\sqrt{b}})\) when \(b>0\)

OpenStudy (zzr0ck3r):

ahh ok

OpenStudy (zzr0ck3r):

\(\frac{1}{x^2+4}=\frac{1}{4}\frac{1}{\frac{x^2}{4}-\frac{4}{4}}=\frac{1}{4}\frac{1}{\frac{x^2}{2^2}-1}=\frac{1}{4}\frac{1}{(\frac{x}{2})^2-1}\)

OpenStudy (zzr0ck3r):

same deal as \(x^2+4=4(\frac{x^2}{4}+1)=4((\frac{x^2}{2})-1)\)

OpenStudy (anonymous):

ooooohh makes sense so when we actor in 1/4 we get what we had before duhh .. wow that was really simple

OpenStudy (zzr0ck3r):

when we factor something out we are really just multiplying by 1 \(a+b=1*(a+b)=\frac{a}{a}(a+b)=a(\frac{a}{a}+\frac{b}{a})=a(1+\frac{b}{a})\)

OpenStudy (zzr0ck3r):

this should also make it clear why we can never "factor out" a zero.:)

OpenStudy (anonymous):

ok yah! dude thanks this is alot of help!

OpenStudy (zzr0ck3r):

np:)